OpenStudy (anonymous):

what is the LCM of 9&15?

OpenStudy (jamesj):

Notice that \( 9 = 3^2 \) and \( 15 = 3^1 \times 5^1 \) hence the LCM is \[ 3^2 \times 5^1 = 45 \]

OpenStudy (anonymous):

thank you.. we are studying for a test. my son has not been doing so well in math and i wanted to double check him. its been a long time since ive done this..

OpenStudy (jamesj):

sure